Starbucks originated in Seattle Washington as a coffee shop. In 2010, there were 10,084 stores
worldwide. By 2014, this had grown to 10,884.
a. Find a linear equation that fits the data x years after 2010.
b. Use the linear equation to predict the number of Starbucks in 2020.
c. In what year is Starbucks predicted to have 18,627 shops worldwide?

Answer by Boreal(15235)

(0, 10084) is the ordered pair for 2010
(4, 10884) is it for 2014
the slope is 800/4=200
point slope formula y-y1=m(x-x1) (x1, y1) point
y-10084=200(x)
y=200x+10084, x is years after 2010.
in 2020, x=10, y=2000+10084 or 12084 stores in 2020.
18627=200x+10084
8543=200x
x=42.715 years after 2010 or about mid to late-2053 (42 years and 8.5 months)